Peter is outside looking at pigs and chickens in the yard. He comes inside, tells Nadia that he counted 13 heads and 36 feet, and asks her how many pigs and chickens were in the yard. what should Nadia tell him?
Found 2 solutions by mananth, texttutoring:
Answer by mananth(16949)

Peter is outside looking at pigs and chickens in the yard. He comes inside, tells Nadia that he counted 13 heads and 36 feet, and asks her how many pigs and chickens were in the yard. what should Nadia tell him?
...
pig =x
chickens=y
x+y=13 heads..............1
4x+2y=36...................2
...
multiply 1 by -2
-2x-2y=-26
add to 2
4x-2x+2y-2y=36-26
2x=10
/2
x= 5 number of pigs
chickens qill be 8
...
m.ananth@hotmail.ca

Answer by texttutoring(324)

You have two variables, number of chickens and number of pigs. This means you need to algebra equations.

Let c = number of chickens
Let p = number of pigs

We know that there are 13 heads in total, so that means that:

c + p = 13

(Because each chicken and pig has one head each!)

We also know that there are 36 feet. Chickens have 2 feet, while pigs have 4 feet:

2c + 4p = 36

Now you have your two equations. Solve them with elimination or substitution.

c+p=13
2c+4p=36

Multiply the first equation by 2, then subtract:
2c+2p=26
2c+4p=36
---------
0 - 2p = -10
p = 5

Now find c:
c = 13 - p
c= 13-5
c = 8

So there are 8 chickens and 5 pigs.